# Prospects for sub-EW ALP searches via $\gamma+b\bar{b}$ signatures at the LHC using jet substructure techniques

### v1

@@ 3.3 Flavour labelling and $b$-tagging on the simulated events
Delphes, used for simulating the detector reconstruction, not having tracking reconstruction capability, uses a flavour dependent parametrized single jet $b$-tagging probability based on ATLAS and CMS published performance ATL-PHYS-PUB-2015-022 ; CMS:2012feb . The emulation of double $b$-tagging is developed and applied outside Delphes by the authors of this work. In particular, we use a parametrisation of the boosted double $b$- and double $c$-tagging algorithm with a working point of $80\%$ $bb$ tagging efficiency. The corresponding rejection efficiencies for multijet background are taken from the GN2X performance listed in Ref. ATL-PHYS-PUB-2023-021 . The parametrisation involves identifying smaller radius jets, called subjets, inside the AK1.0 jet and matching them to a hadron in the event to identify different double flavour jet categories, and it is explained in what follows.
@@
The multijet background events which passes the soft-drop grooming and $\mathrm{N}_{2}^{\mathrm{DDT}}<0$ requirements, discussed Section 3.2, mostly contain AK1.0 jets presenting a two prong substructure. These may originate from the initial state partons in the following three ways: two quarks ($qq^{\prime}$, $q/q^{\prime}=u,d,c,s,b$), or a quark and a gluon ($qg$), or two gluons (gg), produced back-to-back to the high $p_{\rm T}$ photon. In order to categorise the flavour content of these events, ATLAS uses variable radius (VR) anti-$k_{\rm T}$ jets Krohn:2009zg ; ATL-PHYS-PUB-2017-010 , since fixed radius jet algorithmsCacciari:2008gp ; CATANI1993187 ; Wobisch:1998wt introduce overlapping subjets and reduce the efficiency for flavour labelling in case of low ALP mass where the decay products are highly collimated.
Figure 5: Comparison of double subjet $b$-labelling efficiency for different choices of the $\rho$VR parameter while constructing the VR subjets. The efficiencies are shown for the simulated ALP signal with masses, (a) $m_{a}=$ 12 GeV, (b) $m_{a}=$ 40 GeV (c) $m_{a}=$ 60 GeV and (d) $m_{a}=$ 100 GeV. The $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$ values are set to 0.02 and 0.4 ATL-PHYS-PUB-2017-010 , respectively.
The VR algorithm can differentiate the substructure coming from a hard parton than from soft radiation because the radius of the constructed jet changes according to the jet transverse momentum. Three parameters are required to define the VR subjets: $\rho$VR, $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$. $\rho$VR is a dimensionful constant and the effective jet radius depends on the jet transverse momentum as, $R_{\mathrm{eff}}(p_{\rm T})=$$\rho$VR$/p_{\rm T}$. The parameters, $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$, ensure that the effective jet radius do not fall below the detector resolution at high $p_{\rm T}$ and jets do not become arbitrarily large at small $p_{\rm T}$, respectively. These three parameters need to be optimised to improve signal acceptance or background rejection for specific scenarios. The minimum and maximum values for the radius parameter are optimised in Ref. ATL-PHYS-PUB-2017-010 and we use them in our analysis, $R_{\mathrm{min}}$ = 0.02 and $R_{\mathrm{max}}$ = 0.4. However, we optimise $\rho$VR in the following paragraph since a wide range of masses for the ALP signal are considered in this work.
We choose $\rho$VR to maximise the double subjet $b$-labelling efficiency, as a function of the transverse momentum of the AK1.0 jet. The efficiency is defined as the ratio between the number of AK1.0 jets having two $b$-subjets with $p_{\mathrm{T},\text{subjet}}>7$ GeV and the total number of AK1.0 jets. The flavour matching procedure is discussed in the following paragraph  [^3] Note that the same flavour matching method is used first to set the optimal $\rho$VR parameter, and later to categorise the AK1.0 jet into different flavour composition.. A comparison of these efficiencies for different choices of the $\rho$VR parameter is shown in Fig. 5 for ALP masses of 12 GeV, 40 GeV, 60 GeV and 100 GeV. A smaller $\rho$VR value is preferred for a light ALP, while the efficiency is higher with larger $\rho$VR parameter for a heavier ALP. These preferred choices for the $\rho$VR value, specifically $\rho$VR$=5$ GeV for $m_{a}=12$ GeV, 20 GeV and 30 GeV; $\rho$VR$=10$ GeV for $m_{a}=40$ GeV and 50 GeV; $\rho$VR$=20$ GeV for $m_{a}=60$ GeV and 70 GeV; $\rho$VR$=30$ GeV for $m_{a}=80$ GeV, 90 GeV and 100 GeV; are used while constructing the two subjets with the VR algorithm, for both the signal and multijet background.
@@
$p_{\mathrm{T},\text{subjet}}$ $>7$ GeV
@@
$p_{\mathrm{T},\text{hadron}}$ $>5$ GeV
$\Delta R$(hadron, VR) $<0.3$
$\Delta R$(hadron/VR, AK1.0) $<1.0$
Table 3: Summary of the parameters used in variable radius (VR) subjet construction (top panel) and the matching procedure (bottom panel). The radius parameter for the VR subjet is $R_{\mathrm{eff}}(p_{\rm T})=$$\rho$VR$/p_{\rm T}$. The transverse momentum of the VR subjet is $p_{\mathrm{T},\text{subjet}}$. The minimum and maximum radius parameters in the VR algorithm are $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$, respectively. If a $b$-hadron is found, the jet is called $b$-jet, otherwise if a $c$-hadron is found, the jet is called $c$-jet, if no $c$- or $b$-hadrons are present the jet is referred to as light jet. Here, jet refers to the VR subjets.
Since we have at least two flavour matched VR subjets, the AK1.0 jets can be broken down to the following flavour compositions: $bb$, $bc$, $bl$, $cc$, $cl$, and $ll$. The signal events having at least two VR subjets inside the AK1.0 jet labelled as $b$-jets ($bb$) are approximately $71\%$ for a 12 GeV ALP signal, which increases to around $90\%$ for $m_{a}=20$ GeV and 30 GeV. The percentage of signal events passing that criteria decreases gradually afterwards, reducing down to roughly $62\%$ for $m_{a}=100$ GeV, since the jet radius becomes insufficient to include all hadronisation objects, particularly for heavier ALP (see Fig. 5). In the case of the multijet sample, around $85\%$ of the AK1.0 jets have no $b$- or $c$-subjets. The fraction of AK1.0 jets with subjet flavour in the various categories is tabulated in Table 4, for $\rho$VR$=30$ GeV. The fractional values hardly change with $\rho$VR parameter.
@@
Category $bb$ $bc$ $bl$ $ll$ $cl$ $cc$
@@
GN2X efficiency ATL-PHYS-PUB-2023-021 $0.59$ $0.21$ $0.07$ $0.21\times 10^{-3}$ $0.41\times 10^{-2}$ $0.02$
Table 4: The fraction of AK1.0 jets in the multijet sample, sorted by flavour composition, that have at least two VR subjets. The values are shown for $\rho$VR$=30$ GeV and remain roughly the same with different choices of the $\rho$VR parameter. The GN2X efficiency for background in each category are shown for the working point of $80\%$ $bb$ tagging efficiency, from Ref. ATL-PHYS-PUB-2023-021 .
@@ 3.4 Signal extraction and background estimation
The final search is performed by selecting a mass window around the ALP reconstructed $\mathrm{m}_{\mathrm{SD}}$ mass variable after all selection requirements listed in Table 2. For each mass point, we estimate the $\mathrm{m}_{\mathrm{SD}}$ resolution using a double-sided Crystal Ball function Oreglia:1980cs ; ATLAS-CONF-2014-031 , which is found to be robust against the change in shape of the resonance peak for different ALP masses. The Crystal Ball function, consists of a Gaussian function to fit the core of the distrubtion in combination of a power law function to fit both the low and high $\mathrm{m}_{\mathrm{SD}}$ regions. The Crystal Ball function is defined as:
@@
$m_{a}$ [GeV] Fitted parameters [GeV] $\epsilon$ (×$\times 10^{-2}$) Yield at $\mathcal{L}=3000$ fb-1
$\mu$ $\sigma$ $S$ ($g_{\mathrm{aff}}=0.1$ GeV-1 ) $N_{\rm B}$
@@
The mass window, defined as the interval $[\mu-\sigma,\mu+\sigma]$, is used for the search of each ALP mass hypothesis and is tabulated in Table 5, together with the corresponding signal efficiency and multijet background yield for proton-proton collisions at center-of-mass energy $\sqrt{s}=14$ TeV and 3000 fb-1 of integrated luminosity. As discussed earlier, the low signal efficiency at $m_{a}=12$ GeV results from the lack of 2-prong substructure at high jet $p_{\rm T}$ and therefore a low $\mathrm{N}_{2}^{\mathrm{DDT}}$ efficiency. To be noticed that since for high values of the ALP mass, the decay products tend to be less collimated, and therefore less likely to be contained in the cone of the AK1.0 jet, the width of the $\mathrm{m}_{\mathrm{SD}}$ distribution increases.
@@ 4.1 Projected sensitivities on the cross-section and coupling
where $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$ is the signal production cross-section, $\mathcal{L}$ is the integrated luminosity, and $\epsilon$ is the signal efficiency. After repeating the analysis for each $m_{a}$, we obtain $\epsilon(m_{a})$ and $N_{\rm B}(m_{a})$ as listed in Table 5. In Fig. 7, the $95\%$ CL upper limit on $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$ is shown as a function of ALP mass, by the solid blue line in the absence of systematic uncertainty, and by the (smaller) dashed blue line in the case of a $2\%$ ($5\%$) systematic uncertainty  [^4] Adding $x\%$ systematic uncertainty, the signal significance formula is $S/\sqrt{N_{\rm B}+(x\times N_{\rm B}/100)^{2}}$.. The limits are shown for $\sqrt{s}=13$ TeV and $\mathcal{L}=300$ fb-1, in the left panel of Fig. 7, and for $\sqrt{s}=14$ TeV and $\mathcal{L}=3000$ fb-1, in the right panel of Fig. 7. In the latter case, the upper limits are stronger and vary between $6.5\cdot 10^{4}$ pb for an ALP mass of 12 GeV, and 119 pb for an ALP mass of 100 GeV. The limits weaken to $1.1\cdot 10^{5}$ pb ($2.2\cdot 10^{5}$ pb) and 237 pb (524 pb) upon adding $2\%$ ($5\%$) systematic uncertainty, respectively.
@@
Figure 7: Upper limit at $95\%$ C.L. on $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$, as a function of the ALP mass. The limits are shown for (a) $\sqrt{s}=13$ TeV with $\mathcal{L}=300$ fb-1 and (b) $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. The solid and (shorter) dashed blue lines correspond to neglecting systematic uncertainties or including them at the $2\%$ ($5\%$) level, respectively. The solid and dashed gray lines correspond to the theoretical prediction for the cross-section in the ALP model with $g_{\mathrm{aff}}=0.1$ GeV-1 and 0.05 GeV-1, respectively. For comparison, the dashed green line corresponds to the projected upper limit on the cross-section for an analogous process, $\sigma(pp\to Z^{\prime}\gamma\to q\bar{q}\gamma)$ from Ref. CMS:2019xai .
Figure 8: Upper limit at $95\%$ C.L. on the ALP-fermion coupling $g_{\mathrm{aff}}$, as a function of the ALP mass. The limits are shown for (a) $\sqrt{s}=13$ TeV with $\mathcal{L}=300$ fb-1 and (b) $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. The solid and (shorter) dashed blue lines correspond to neglecting systematic uncertainties or including them at the $2\%$ ($5\%$) level, respectively. The dashed green and brown lines correspond to the projected upper limits from Ref. CMS:2019xai and Ref. CMS:2017dcz , respectively.
The derived upper limits on the signal production cross-section are translated into upper limits on the ALP-fermion coupling $g_{\mathrm{aff}}$, as illustrated as a function of $m_{a}$ in Fig. 8. The left and right panel corresponds to upper limits at $\sqrt{s}=13$ TeV with $\mathcal{L}=300$ fb-1 and $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1, respectively. The couplings above the blue solid and (shorter) dashed line can be excluded at $95\%$ C.L. assuming zero or $2\%$ ($5\%$) systematic uncertainty, respectively. We observe that the stronger upper limit is obtained for an ALP mass $m_{a}\simeq 40$ GeV, with $g_{\mathrm{aff}}\leq 0.053$ GeV-1 at $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. For larger masses the limit slightly weakens, varying between $g_{\mathrm{aff}}\lesssim(0.056$ $0.098$) GeV-1 in the range $m_{a}\sim(50-100)$ GeV. Adding $2\%$ and $5\%$ systematic uncertainty, the coupling reach degrades to $g_{\mathrm{aff}}\lesssim(0.076$ $0.27$) GeV-1 and (0.11(0.11( 0.11 - $0.39$) GeV-1, respectively, in the ALP mass range between 10 GeV and 100 GeV. We compare these limits with the results available at the current LHC run at $\sqrt{s}=13$ TeV with $\mathcal{L}=35.9$ fb-1: the dashed green and brown lines are the projected $95\%$ C.L. upper limits on the coupling to SM quarks, in the searches for a $Z^{\prime}$ resonance in association with a photon CMS:2019xai or a jet CMS:2017dcz , respectively. We observe that our limits are stronger as compared to these references, where the analyses were performed without any flavour decomposition of the jet.

### v3

@@ 3.3 Flavour labelling and $b$-tagging on the simulated events
Delphes, used for simulating the detector reconstruction, not having tracking reconstruction capability, uses a flavour dependent parametrized single jet $b$-tagging probability based on ATLAS and CMS published performance ATL-PHYS-PUB-2015-022 ; CMS:2012feb . The emulation of double $b$-tagging is developed and applied outside Delphes by the authors of this work. In particular, we use a parametrisation of the boosted double $b$- and double $c$-tagging algorithm with a working point of $80\%$ bbbb tagging efficiency. The corresponding rejection efficiencies for multijet background are taken from the GN2X performance listed in Ref. ATL-PHYS-PUB-2023-021 . The parametrisation involves identifying smaller radius jets, called subjets, inside the AK1.0 jet and matching them to a hadron in the event to identify different double flavour jet categories, and it is explained in what follows.
@@
The multijet background events which passes the soft-drop grooming and $\mathrm{N}_{2}^{\mathrm{DDT}}<0$ requirements, discussed Section 3.2, mostly contain AK1.0 jets presenting a two prong substructure. These may originate from the initial state partons in the following three ways: two quarks ($q^{\prime}, q/q$$q^{\prime}=u,d,c,s,b$), or a quark and a gluon ($qg$), or two gluons (gg), produced back-to-back to the high $p_{\rm T}$ photon. In order to categorise the flavour content of these events, ATLAS uses variable radius (VR) anti-$k_{\rm T}$ jets Krohn:2009zg ; ATL-PHYS-PUB-2017-010 , since fixed radius jet algorithmsCacciari:2008gp ; CATANI1993187 ; Wobisch:1998wt introduce overlapping subjets and reduce the efficiency for flavour labelling in case of low ALP mass where the decay products are highly collimated.
@@
[…]
Next, we match each VR subjet to a hadron. To be compliant with the labelling procedure used to derive the parametrized efficiency in Ref. ATL-PHYS-PUB-2020-019 , we assign a $b$-, $c$- and light flavor, to each VR subjet. We first identify $b$-hadrons and $c$-hadrons in the event from the list of all stable particles provided by Pythia8 using their PDG codes. The hadrons must have $p_{\mathrm{T},\text{hadron}}>5$ GeV, and are required to be within $\Delta R<1.0$ of the AK1.0 jet axis. Events are selected if they contain at least two VR subjets with $p_{\mathrm{T},\text{subjet}}>7$ GeV, similar to earlier double subjet $b$-labelling efficiency, and the subjets fall inside the AK1.0 jet. Thus the selected events have at least two VR subjets and identified hadrons within $\Delta R<1.0$ of the AK1.0 jet. The VR subjet is called $b$-jet (b) when a $b$-hadron is found within $\Delta R<0.3$ ATL-PHYS-PUB-2020-019 of the VR subjet. If there is no nearby $b$-hadron but a $c$-hadron is found with $\Delta R(\text{$c$-hadron, VR-subjet})<0.3$, the VR subjet is labelled as $c$-jet (c). In case the $b$-hadron or $c$-hadron is within $\Delta R<0.3$ of multiple VR subjets, the hadron is associated to the closest VR subjet. Otherwise, the VR subjet is labelled as a light jet (l). We summmarise all the above discussion in Table 3.
@@
$p_{\mathrm{T},\text{subjet}}$ >7>7 GeV
@@
$p_{\mathrm{T},\text{hadron}}$ >5>5 GeV
$\Delta R$(hadron, VR) <0.3<0.3
$\Delta R$(hadron/VR, AK1.0) <1.0<1.0
[…]
@@
Category bbbb bcbc blbl llll clcl cccc
Fractional composition ($\%$) $0.6$ $0.5$ $1.2$ $84.4$ $10.5$ $2.8$
GN2X efficiency ATL-PHYS-PUB-2023-021 $0.59$ $0.21$ $0.07$ $0.21\times 10^{-3}$ $41\times 10^{-2} 0.0$20.02
@@ 3.4 Signal extraction and background estimation
In light of the precise background estimation provided by the $\mathrm{m}_{\mathrm{SD}}$ side-band fit, the final result is derived by counting the number of signal and background events, after the selection requirements listed in Table 2, in a mass window centered around the ALP reconstructed $\mathrm{m}_{\mathrm{SD}}$. While complex analysis strategies involving substructure observables could increase the sensitivity of the resonance search and should be attempted in the experimental analysis, these are not considered in this work since such approaches would require several steps that use collider data for validation. For each mass point, we estimate the $\mathrm{m}_{\mathrm{SD}}$ resolution using a double-sided Crystal Ball function Oreglia:1980cs ; ATLAS-CONF-2014-031 , which is found to be robust against the change in shape of the resonance peak for different ALP masses. The Crystal Ball function, consists of a Gaussian function to fit the core of the distribution in combination of a power law function to fit both the low and high $\mathrm{m}_{\mathrm{SD}}$ regions. The Crystal Ball function is defined as:
@@
$m_{a}$ [GeV] Fitted parameters [GeV] $\epsilon$ ($\times 10^{-2}$) Yield at $\mathcal{L}=3000$ fb-1
$\mu$ $\sigma$
$S$ ($g_{\mathrm{aff}}=0.1$ GeV-1
$N_{\rm B}$
@@
The mass window, defined as the interval $\mu-\sigma,\mu+\sigma]$, is used for the search of each ALP mass hypothesis and is tabulated in Table 5, together with the corresponding signal efficiency and multijet background yield for proton-proton collisions at center-of-mass energy $\sqrt{s}=14$ TeV and 3000 fb-1 of integrated luminosity. In this work, the multijet background is estimated with the generators chosen in line with what is used by LHC collaborations ATL-PHYS-PUB-2021-028 ; ATLAS:2018hbc ; CMS:2019xai ; CMS:2017dcz . The real experimental analysis will have to implement data-driven background techniques, which on one side will render the analysis insensitive to modeling uncertainties, and on the other side will allow the reduction of statistical uncertainties. This is due to the large sample of QCD events recorded by the collaboration, whose size is hard to match by simulations, given the amount of resources necessary to generate a comparable number of multijet events.
As discussed earlier, the low signal efficiency at $m_{a}=12$ GeV results from the lack of 2-prong substructure at high jet $p_{\rm T}$ and therefore a low $\mathrm{N}_{2}^{\mathrm{DDT}}$ efficiency. To be noticed that since for high values of the ALP mass, the decay products tend to be less collimated, and therefore less likely to be contained in the cone of the AK1.0 jet, the width of the $\mathrm{m}_{\mathrm{SD}}$ distribution increases.
@@ 4.1 Projected sensitivities on the cross-section and coupling
where $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$ is the signal production cross-section, $\mathcal{L}$ is the integrated luminosity, and $\epsilon$ is the signal efficiency. After repeating the analysis for each $m_{a}$, we obtain $\epsilon(m_{a})$ and $N_{\rm B}(m_{a})$ as listed in Table 5. In Fig. 7, the $95\%$ CL upper limit on $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$ is shown as a function of ALP mass, by the solid blue line in the absence of systematic uncertainty, and by the (smaller) dashed blue line in the case of a $2\%$ ($5\%$) systematic uncertainty  [^4] Adding $x\%$ systematic uncertainty, the signal significance formula is $S/\sqrt{N_{\rm B}+(x\times N_{\rm B}/100)^{2}}$.. The size of the systematic uncertainties chosen are in line with current estimations by LHC experiments for jet mass scale and resolution CMS:2019xai . The limits are shown for $\sqrt{s}=13$ TeV and $\mathcal{L}=35.9$ fb-1, in the left panel of Fig. 7, and for $\sqrt{s}=14$ TeV and $\mathcal{L}=3000$ fb-1, in the right panel of Fig. 7. In the latter case, the upper limits are stronger and vary between $6.5\cdot 10^{4}$ pb for an ALP mass of 12 GeV, and 119 pb for an ALP mass of 100 GeV. The limits weaken to $1.1\cdot 10^{5}$ pb ($2.2\cdot 10^{5}$ pb) and 237 pb (524 pb) upon adding $2\%$ ($5\%$) systematic uncertainty, respectively.
@@
Figure 7: Upper limit at $95\%$ C.L. on $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$, as a function of the ALP mass. The limits are shown for (a) $\sqrt{s}=13$ TeV with $\mathcal{L}=35.9$ fb-1 and (b) $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. The solid and (shorter) dashed blue lines correspond to neglecting systematic uncertainties or including them at the $2\%$ ($5\%$) level, respectively. The solid and dashed gray lines correspond to the theoretical prediction for the cross-section in the ALP model with $g_{\mathrm{aff}}=0.1$ GeV-1 and 0.05 GeV-1, respectively. For comparison, the dashed green line corresponds to the projected upper limit on the cross-section for an analogous process, $\sigma(pp\to Z^{\prime}\gamma\to q\bar{q}\gamma)$ from Ref. CMS:2019xai .
Figure 8: Upper limit at $95\%$ C.L. on the ALP-fermion coupling $g_{\mathrm{aff}}$, as a function of the ALP mass. The limits are shown for (a) $\sqrt{s}=13$ TeV with $\mathcal{L}=35.9$ fb-1 and (b) $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. The solid and (shorter) dashed blue lines correspond to neglecting systematic uncertainties or including them at the $2\%$ ($5\%$) level, respectively. The dashed green line correspond to the projected upper limits from Ref. CMS:2019xai
The derived upper limits on the signal production cross-section are translated into upper limits on the ALP-fermion coupling $g_{\mathrm{aff}}$, as illustrated as a function of $m_{a}$ in Fig. 8. The left and right panel corresponds to upper limits at $\sqrt{s}=13$ TeV with $\mathcal{L}=35.9$ fb-1 and $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1, respectively. The couplings above the blue solid and (shorter) dashed line can be excluded at $95\%$ C.L. assuming zero or $2\%$ ($5\%$) systematic uncertainty, respectively. We observe that the stronger upper limit is obtained for an ALP mass $m_{a}\simeq 60$ GeV, with $g_{\mathrm{aff}}\leq 0.045$ GeV-1 at $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. For larger masses the limit slightly weakens, varying between $g_{\mathrm{aff}}\lesssim(0.047$ - $0.060$) GeV-1 in the range $m_{a}\sim(70-100)$ GeV. Adding $2\%$ and $5\%$ systematic uncertainty, the coupling reach degrades to $g_{\mathrm{aff}}\lesssim(0.084$ - $0.26$) GeV-1 and $(0.12$ - $0.38$) GeV-1, respectively, in the ALP mass range between 12 GeV and 100 GeV. We compare these limits with the results available at the current LHC run at $\sqrt{s}=13$ TeV with $\mathcal{L}=35.9$ fb-1: the dashed green line is the projected $95\%$ C.L. upper limits on the coupling to SM quarks, in the searches for a $Z^{\prime}$ resonance in association with a photon CMS:2019xai . We observe that our limits are stronger as compared to this reference, where the analyses were performed without any flavour decomposition of the jet.